OpenStudy (anonymous):

Simplify 5√20 - 2√45

OpenStudy (anonymous):

\[5\sqrt{2*2*5} - 2\sqrt{3*3*5} = 5*2*\sqrt{5} - 2*3*\sqrt{5}\] \[10\sqrt{5} - 6\sqrt{5} = 4\sqrt{5}\]

OpenStudy (anonymous):

Hey kvhu, if u r satisfied with the answer, pls become my fan as i need just one more to complete 50 !!!!! :})